# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

### v1

@@ VI-B The Case where Only System Behaviors Occur
The case where only system behaviors occur, i.e., the optimal trajectory consists of a finite number of unconstrained arcs and constrained arcs that are not tangent to the constrained boundary, is widely applied in existing works on trajectory planning [31, 32, 8].
Denote the augmented switching law of the trajectory as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}{\mathcal{E}}_{N}$, where ${\mathcal{E}}_{N}$ is induced by ${\boldsymbol{x}}\left(t_{N}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$. $\forall k\in{\mathbb{N}}$, denote $\left|\overline{k}\right|=\left|\underline{k}\right|=k$, ${\mathrm{sgn}}\left(\overline{k}\right)=+1$, and ${\mathrm{sgn}}\left(\underline{k}\right)=-1$. The non-existence of tangent markers and additional end-constraints in ${\mathcal{L}}$ means that $\forall i\in\left[N\right]$, the arc ${\mathcal{S}}_{i}$ achieves strictly feasibility. An arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ in $\left[t_{0},t_{1}\right]$ is strictly feasible if $\forall p\not\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ holds in $\left[t_{0},t_{1}\right]$.
@@
Assume that the augmented switching law consists of system behaviors without tangent markers and additional end-constraints except at ${t_{\mathrm{f}}}$. Assume that:
@@
By Condition 3, $\left|{\mathcal{S}}_{N}\right|=0$. Denote $\left|{\mathcal{S}}_{N+1}\right|=n$, and =$\mathcal{I}=\left\{i\in\left[N+1\right]:\left|{\mathcal{S}}_{i}\right|\not=0\right\}$. According to Corollary 1, ${\boldsymbol{J}}$ contains the following rows:
@@
Let ${\boldsymbol{A}}\leftrightarrow{\boldsymbol{B}}$ mean that ${\boldsymbol{A}}$ has full row rank if and only if ${\boldsymbol{B}}$ has full row rank. Since $\forall j\in\left[N-1\right]$, $\Delta u_{j+1}\not=0$, it has
@@
Further research suggests that Conditions 1, 2, and 3 in Corollary 3 might hold almost everywhere. In other words, $\left({\boldsymbol{x}}_{0},{\boldsymbol{x}}_{\mathrm{f}},{\boldsymbol{x}}_{\mathrm{m}},{u_{\mathrm{m}}}\right)$ whose resulting Bang-Singular-Bang trajectory does not satisfy Conditions 1, 2, and 3 forms a zero-measure set.
Remark.
In our previous work [8], a similar conclusion pointed out that when fixing ${{\boldsymbol{x}}_{\mathrm{f}}}$, then the set of $\left({\boldsymbol{x}}_{0},{\boldsymbol{t}}\right)$ that satisfies the augmented switching law ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}{\mathcal{E}}_{N}$ locally forms a submanifold of $r$-dimensions. Among them, $r=N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|$. However, the local optimality is not discussed in [8]. Corollary 3 proves that the augmented switching law of dimensions $r>n$ fails to achieve optimality.
${\boldsymbol{J}}=\left[\begin{array}[]{ccccccccc}-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{3}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3}-t_{2}\right)&0&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{6}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{6}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{5}\right)&0&\dots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&\dots&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}\\ -2{u_{\mathrm{m}}}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{5}\right)&0&\dots&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-1}\right)\\ -2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{2}\right)&{\boldsymbol{0}}&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{5}\right)&{\boldsymbol{0}}&\dots&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-1}\right)\\ \end{array}\right]$
$\left[\begin{array}[]{cc}\left(\delta_{j=i}-\delta_{j=i-1}\right)_{i\in\left[N\right],j\in\left[N\right]}&\\ &{\boldsymbol{I}}_{n-1}\end{array}\right]\left[\begin{array}[]{cccc}{\boldsymbol{I}}_{N-1}&&&\\ &0&1&\\ &1&0&\\ &&&{\boldsymbol{I}}_{n-2}\end{array}\right]\left[\begin{array}[]{cc}{\boldsymbol{I}}_{N+1}&\\ &\left(\phi_{i+2-j}\left(t_{\infty}-t_{3N}\right)\right)_{i\in\left[n-2\right],j\in\left[n\right]}\end{array}\right]$ (63)
It is challenging to reason Corollary 3 by existing costate-based necessary conditions [14] due to the complex behavior of the costates. In contrast, existing costate-based necessary conditions can try to reason the costate for a given trajectory case-by-case.
For example, in a 4th order problem, $\mathcal{L}_{1}=\underline{01}\overline{01}\underline{0}\overline{2}\underline{01}\overline{01}\underline{0}$ can be feasible, but it cannot be optimal since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=5>4$. $\mathcal{L}_{2}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$ can serve as a candidate for the optimal control since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=4$. Of course, both $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ cannot be optimal in a 3rd order problem. Though Corollary 3 is not a sufficient condition for optimal control, it provides a simple and effective approach to analyzing the optimal control of problem (53a).
@@
Our previous work [18] points out that if the chattering phenomenon occurs in problem (53a), then there exists one and only one inequality state constraint that switches between active and inactive during the chattering phenomenon. An infinite number of unconstrained arcs are joined by the constrained boundary, i.e., the unconstrained arcs are tangent to the boundary for infinite times. In contrast, constrained arcs do not occur. This section discusses the chattering phenomenon induced by $x_{2}\leq x_{{\mathrm{m}}2}$ based on the proposed state-centric necessary condition. [18] provides the following conclusion.
@@
Among them, $t_{3k-1}=\frac{3t_{3k}+t_{3k-3}}{4}$ and $t_{3k-2}=\frac{t_{3k}+3t_{3k-3}}{4}$. Hence, $\Delta u_{3k-1}=2{u_{\mathrm{m}}}$, $\Delta u_{3k}=-2{u_{\mathrm{m}}}$, and $\Delta u_{3k+1}=0$.
$\forall N\in{\mathbb{N}}^{*}$, the augmented switching law between ${\boldsymbol{x}}\left(t_{0}\right)$ and ${\boldsymbol{x}}\left(t_{3N}\right)$ is denoted as ${\mathcal{L}}_{N}$, which requires that $\forall k\in\left[N-1\right]$, ${\boldsymbol{x}}_{3k,1:2}=x_{{\mathrm{m}}2}{\boldsymbol{e}}_{2}$, and ${\boldsymbol{x}}_{3N}={\boldsymbol{x}}\left(t_{3N}\right)$. According to Corollary 1, the Jacobian matrix ${\boldsymbol{J}}$, except the last column, is (62). $\forall k\in\left[N-1\right]$, the $3k$-th columns and the $\left(2k-1\right)$-th rows can be eliminated, and the coefficients of each column, i.e., ±$\pm 2{u_{\mathrm{m}}}$, can be normalized, resulting in ${\boldsymbol{J}}\leftrightarrow{\boldsymbol{J}}_{1}$. Upon left multiplication by (63) and right multiplication by $\left(\delta_{j=i}-\delta_{j=i+1}\right)_{i\in\left[N\right],j\in\left[N\right]}^{3}$, it has ${\boldsymbol{J}}\leftrightarrow{\boldsymbol{J}}^{\prime}$. Considering the rank of columns $\left(k-n+3\right)$ through $k$ of ${\boldsymbol{J}}^{\prime}$, (66) holds. ∎
@@
According to Corollary 3, the chattering phenomenon can be induced by $x_{2}\leq x_{{\mathrm{m}}2}$ only if $\lim_{k\to\infty}\tau_{k}=0$ and $\tau_{k}>0$ when applying (66). When given initial values, (66) is a polynomial equation w.r.t. $\tau_{k}$. It is evident that $n=3$ does not hold since it has $\tau_{k}=2\tau_{k-1}-\tau_{k-2}$. Through more refined derivation, it can be rigorously proven that $n=4$ does not hold. The proof is lengthy and omitted due to space limitation, whose approach is similar to our previous work [18]. In other words, the chattering phenomenon induced by $x_{2}\leq x_{{\mathrm{m}}2}$ can only occur in a 5th order problem or higher. One can try to determine whether the chattering phenomenon can occur in an $n$-th order problem for $n\geq 5$ based on the proposed Corollary 3.
Remark.
@@
It is challenging to obtain the recursive equation (66) for problem (53a) of high order based on the traditional costate-based necessary condition [14] since the behaviors of costates are significantly complex for high order problems.
@@
Although Theorem 3 holds under the assumption of the non-existence of the chattering phenomenon, Corollary 4 provides an example to apply the proposed necessary condition to the case the chattering phenomenon occurs, where a sub-arc without chattering of the optimal trajectory is investigated.
VI-D Simulation Experiment
Figure 2: Results of the simulation experiment. (a) The plot of the local minimal $t_{\mathrm{f}}^{\prime}-t_{\mathrm{f}}$ and $t_{4}^{\prime}-t_{4}$. (b) The original trajectory and the optimized trajectory. Among them, ${\boldsymbol{x}}_{0}=\left(0.75,-0.375,2,9\right)$, ${{\boldsymbol{x}}_{\mathrm{f}}}=\left(0.25,0.5,-2,-5\right)$, ${\boldsymbol{x}}_{\mathrm{m}}=\left(1,1.5,4,20\right)$, ${u_{\mathrm{m}}}=1$.

### v2

@@ VI-B The Case where Only System Behaviors Occur
The case where only system behaviors occur, i.e., the optimal trajectory consists of a finite number of unconstrained arcs and constrained arcs that are not tangent to constraints’ boundaries, is widely applied in existing works on trajectory planning [19, 34, 20].
Denote the ASL of the trajectory as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}{\mathcal{E}}_{N}$, where ${\mathcal{E}}_{N}$ is induced by ${\boldsymbol{x}}\left(t_{N}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$. $\forall k\in{\mathbb{N}}$, denote $\left|\overline{k}\right|=\left|\underline{k}\right|=k$, ${\mathrm{sgn}}\left(\overline{k}\right)=+1$, and ${\mathrm{sgn}}\left(\underline{k}\right)=-1$. The non-existence of tangent markers and additional end-constraints in ${\mathcal{L}}$ means that $\forall i\in\left[N\right]$, the arc ${\mathcal{S}}_{i}$ achieves strictly feasibility. An arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ in $\left[t_{0},t_{1}\right]$ is strictly feasible if $\forall p\not\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ holds in $\left[t_{0},t_{1}\right]$.
@@
Assume that the ASL consists of system behaviors without tangent markers and additional end-constraints except at ${t_{\mathrm{f}}}$. Assume that:
@@
$\left[\begin{array}[]{ccccccccc}-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{3}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3}-t_{2}\right)&0&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{6}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{6}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{5}\right)&0&\dots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&\dots&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}\\ -2{u_{\mathrm{m}}}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{5}\right)&0&\dots&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-1}\right)\\ -2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{2}\right)&{\boldsymbol{0}}&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{5}\right)&{\boldsymbol{0}}&\dots&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-1}\right)\\ \end{array}\right]$
@@
By Condition 3, $\left|{\mathcal{S}}_{N}\right|=0$. Denote $\left|{\mathcal{S}}_{N+1}\right|=n$, and =$\mathcal{I}=\left\{i\in\left[N+1\right]:\left|{\mathcal{S}}_{i}\right|\not=0\right\}$. According to Proposition 8, ${\boldsymbol{J}}$ contains the following rows:
@@
Let ${\boldsymbol{A}}\leftrightarrow{\boldsymbol{B}}$ mean that ${\boldsymbol{A}}$ has full row rank if and only if ${\boldsymbol{B}}$ has full row rank. By $\forall j\in\left[N-1\right]$, $\Delta u_{j+1}\not=0$, it holds that
@@
In our previous work [20], a similar conclusion pointed out that when fixing ${{\boldsymbol{x}}_{\mathrm{f}}}$, then the set of $\left({\boldsymbol{x}}_{0},{\boldsymbol{t}}\right)$ that satisfies the ASL ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$ locally forms a submanifold of dimension $r$. Among them, $r=N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|$. However, the local optimality is not discussed in [20]. Corollary 2 proves that the ASL with $r>n$ fails to achieve optimality.
It is challenging to reason Corollary 2 by existing costate-based necessary conditions [9] due to complex behaviors of the costates. In contrast to Corollary 2, PMP-based approaches need to derive the costates for a given trajectory case-by-case.
For example, in a 4th-order problem, $\mathcal{L}_{1}=\underline{01}\overline{01}\underline{0}\overline{2}\underline{01}\overline{01}\underline{0}$ can be feasible, but it cannot be optimal since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=5>4$. $\mathcal{L}_{2}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$ can be a candidate for optimal control since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=4$. Though Corollary 2 is not a sufficient condition for optimal control, it provides a simple and effective approach to analyzing optimal control of problem (48a). Numerical examples are provided in Section VII.
@@
Our previous work [32] points out that if the chattering phenomenon occurs in problem (48a), then there exists one and only one inequality state constraint that switches between active and inactive during the chattering phenomenon. An infinite number of unconstrained arcs are joined by constraints’ boundaries, i.e., the unconstrained arcs are tangent to the boundary for infinite times. In contrast, constrained arcs do not occur. This section discusses the chattering phenomenon induced by $x_{2}\leq x_{{\mathrm{m}}2}$ based on the proposed state-centric necessary condition. [32] provides the following lemma.
@@
Among them, $t_{3k-1}=\frac{3t_{3k}+t_{3k-3}}{4}$ and $t_{3k-2}=\frac{t_{3k}+3t_{3k-3}}{4}$.
$\forall N\in{\mathbb{N}}^{*}$, the ASL between ${\boldsymbol{x}}\left(t_{0}\right)$ and ${\boldsymbol{x}}\left(t_{3N}\right)$ is denoted as ${\mathcal{L}}_{N}$, which requires that $\forall k\in\left[N-1\right]$, ${\boldsymbol{x}}_{3k,1:2}=x_{{\mathrm{m}}2}{\boldsymbol{e}}_{2}$, and ${\boldsymbol{x}}_{3N}={\boldsymbol{x}}\left(t_{3N}\right)$. By Proposition 8, the Jacobian matrix ${\boldsymbol{J}}$, except the last column, is (51). Through basic row and column transformations, it holds that ${\boldsymbol{J}}\leftrightarrow{\boldsymbol{J}}^{\prime}$. Considering the rank of columns $\left(k-n+3\right)$ through $k$ of ${\boldsymbol{J}}^{\prime}$, (59) holds.
@@
According to Corollary 3, the chattering phenomenon can be induced by $x_{2}\leq x_{{\mathrm{m}}2}$ only if $\lim_{k\to\infty}\tau_{k}=0$ and $\tau_{k}>0$ under the recursive equation (59). It is evident that chattering does not occur when $n=3$ since $\tau_{k}=2\tau_{k-1}-\tau_{k-2}$ converges to $\infty$. Through more refined derivation, it can be rigorously proved that $n=4$ does not hold. In future works, one can try to determine the existence of chattering in problem (48a) of order $n\geq 5$ under constraints on $x_{2}$, where Corollary 3 and Theorem 3 would serve as useful mathematical tools.
@@
However, it is challenging to obtain the recursive equation (59) for problem (48a) based on the traditional PMP-based necessary condition [9] since the behaviors of costates are significantly complex for high-order problems, let alone determining the existence of chattering. A comparison between state-centric and PMP-based approaches is provided in Appendix A.
@@
Although Theorem 3 holds under the assumption of the non-existence of the chattering phenomenon, Corollary 3 provides an example to apply the proposed necessary condition to a chattering trajectory, where a sub-arc without chattering of the optimal trajectory is investigated.
VII Numerical Experiments
Numerical experiments for high-order chain-of-integrator systems (48a) are conducted to provide a potential way to apply the proposed state-centric necessary condition to optimizing trajectories. Since this paper focuses on the theoretical necessary condition instead of optimization algorithms, metrics like computational time are not considered. Note that this paper does not provide a manual optimization algorithm.
VII-A Representating a Feasible BBS Trajectory
Consider problem (1a) whose system matrix is non-diagonalizable with multiple Jordan blocks as follows:
${\boldsymbol{A}}=\left[\begin{array}[]{cccc}-1&1&0&0\\ 0&-1&1&0\\ 0&0&-1&0\\ 0&0&0&0\end{array}\right],\,{\boldsymbol{b}}=\left[\begin{array}[]{c}0\\ 0\\ 2\\ -1\end{array}\right].$ (61)
The constraints are $\left|u\right|\leq 1$, $x_{1}\geq-0.7$, and $x_{3}+x_{4}\leq 0.5$. As defined in Proposition 2, $x_{1}\geq-0.7$ is a 3rd-order state constraint. As shown in Figs. 4(a-c), a feasible BBS trajectory ${\boldsymbol{x}}_{\text{ASL}}$ consists of full features defined in this paper. The ASL is ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}{\mathcal{S}}_{3}{\mathcal{S}}_{4}\left({\mathcal{S}}_{5},\left\{{\mathcal{T}}_{5}^{(1)}\right\}\right){\mathcal{S}}_{6}{\mathcal{E}}_{6}{\mathcal{S}}_{7}{\mathcal{S}}_{8}$. Specifically, ${\mathcal{S}}_{1}$, ${\mathcal{S}}_{4}$, and ${\mathcal{S}}_{8}$ are unconstrained arcs where $u\equiv-1$. ${\mathcal{S}}_{2}$, ${\mathcal{S}}_{5}$, and ${\mathcal{S}}_{7}$ are unconstrained arcs where $u\equiv+1$. ${\mathcal{S}}_{3}$ and ${\mathcal{S}}_{6}$ are constrained arcs where $x_{3}+x_{4}\equiv 0.5$. ${\mathcal{T}}_{5}^{(1)}$ is a tangent markers where $x_{1}$ is tangent to $-0.7$. ${\mathcal{E}}_{6}$ is an additional end-constraint requiring $u\left(t_{6}^{-}\right)=1$. In other words, the control $u$ is continuous at the connection of the constrained arc ${\mathcal{S}}_{6}$ and the unconstrained arc ${\mathcal{S}}_{7}$. In contrast, $u$ is not continuous at the two ends of another constrained arc ${\mathcal{S}}_{3}$. One can check that Theorem 3 holds for ${\boldsymbol{x}}_{\text{ASL}}$.
A numerical optimal control toolbox, i.e., Yop [35], is applied to solving problem (61). The number of intervals is set to 200. The resulting trajectory without a specified initial solution ${\boldsymbol{x}}_{\text{Yop1}}$ is significantly far from optimal, as shown in Fig. 4. If setting ${\boldsymbol{x}}_{\text{ASL}}$ as the initial solution, then a near-optimal solution ${\boldsymbol{x}}_{\text{Yop2}}$ is obtained. The terminal times of ${\boldsymbol{x}}_{\text{ASL}}$, ${\boldsymbol{x}}_{\text{Yop1}}$, and ${\boldsymbol{x}}_{\text{Yop2}}$ are 5.541, 6.896, and 5.624, respectively.
Furthermore, costates can be reconstructed from states ${\boldsymbol{x}}_{\text{ASL}}$ through meticulous derivations, as shown in Fig. 4(d). The information on costates also verifies the proposed state-centric necessary condition in a way, noting that PMP (7) holds in this example. In contrast, the costates of ${\boldsymbol{x}}_{\text{Yop1}}$ and ${\boldsymbol{x}}_{\text{Yop2}}$ do not exist since Yop does not solve a problem based on costates.
Figure 4: A feasible BBS trajectory with system behaviors, tangent markers, and additional end-constraints. In (d), $\hat{\lambda}\triangleq\left(2\lambda_{3}-\lambda_{4}\right)\xi\left(t\right)$ where $\xi\left(t\right)>0$ serves as a scaling factor to enhance the visibility of ${\mathrm{sgn}}\left(2\lambda_{3}-\lambda_{4}\right)$.
Figure 5: Optimizing a 4th-order near-optimal trajectory. (a) Plot of the locally minimal $t_{\mathrm{f}}^{\prime}-t_{\mathrm{f}}$ and $t_{4}^{\prime}-t_{4}$. (b) The original and the optimized trajectories.
VII-B Determining the Optimality of a 4th-Order Trajectory